### v1

@@ 5 Conclusion and Outlook
We have demonstrated the operation of a drift-stable, magnetically silent magnetometer with a sensitivity of 3.9 pT and a drift stability of 90 fT over 150 seconds. This meets the requirements of the TUCAN nEDM experiment for magnetometry [25]. Further studies demonstrating the operation of multiple sensors, measuring crosstalk, offset, and light shift are being planned for the completion of the TUCAN magnetically shielded room, currently under construction at TRIUMF. The final magnetometric array will have 20 sensors serviced by the same probe laser, eliminating the effects of laser frequency and intensity variation on differential field measurements.
@@ Acknowledgements
The authors thank Martin Rosner and Georg Bison for helpful communications during development. The authors also thank all of our TUCAN colleagues for support and guidance through these studies. W. Klassen wishes to thank Dr. Derek Fujimoto for help copy-editing the final draft.
@@
This research was supported by the Canada Foundation for Innovation; the Canada Research Chairs program; the Natural Sciences and Engineering Research Council of Canada (NSERC) SAPPJ-2016-00024, SAPPJ-2019-00031 and SAPPJ-2023-00029; Japan Society for the Promotion of Science (JSPS) KAKENHI grants 18H05230 and 20KK0069; JSPS KENHI grants 17K14307, 19K23442, 20K14487, 21K13940, 22H01236; RCNP CORENet; and the KEK-TRIUMF Exchange Program for Early Career Researchers (EPECR). Chris Hovde and Southwest Sciences were contracted for the design and construction of the fiberized sensors.

### v2

@@ 5 Conclusion and Outlook
We have demonstrated the simultaneous operation of multiple drift-stable, magnetically silent magnetometers with a sensitivity of 3.5 pT/$\sqrt{\mathrm{Hz}}$ at 1 Hz and a drift stability of 90 fT over 150 seconds. This meets the requirements of the TUCAN nEDM experiment for magnetometry [25]. Further studies demonstrating the operation of multiple sensors, measuring crosstalk, offset, and light shift are being planned for the completion of the TUCAN magnetically shielded room, currently under construction at TRIUMF. The final magnetometric array will have 20 sensors serviced by the same probe laser, eliminating the effects of laser frequency and intensity variation on differential field measurements.
@@ Acknowledgements
The authors thank Drs. Martin Rosner and Georg Bison for helpful communications during development. The authors also thank all of our TUCAN colleagues for support and guidance through these studies. W. Klassen wishes to thank Dr. Derek Fujimoto for help copy-editing the final draft.
@@
This research was supported by the Canada Foundation for Innovation; the Canada Research Chairs program; the Natural Sciences and Engineering Research Council of Canada (NSERC) SAPPJ-2016-00024, SAPPJ-2019-00031 and SAPPJ-2023-00029. This work was additionally supported by JSPS KAKENHI (Grant Nos. 18H05230, 20KK0069, 20K14487 and 22H01236), JSPS Bilateral Program (Grant No. JSPSBP120239940), JST FOREST Program (Grant No. JPMJFR2237), International Joint Research Promotion Program of Osaka University, RCNP COREnet, the Yamada Science Foundation and the Murata Science Foundation. Chris Hovde and Southwest Sciences were contracted for the design and construction of the fiberized sensors.